A.) A bullet is fired from a gun perfectly horizontal towards the center of a target 200 m away. If the bullet strikes 10 cm below the target, what was the velocity of the bullet?
B.) If the same gun is fired a second time, from the same distance, aimed at the center of the target, but this time strikes 10 cm below and 8 cm to the right of the center of the target, what was the velocity of the wind?

Answer by jorel1380(3719)

A.)The equation for the height of a falling object is h(t)=-4.9 tē+vt+c, where t is time, in seconds, h is height in meters, v is vertical velocity, and c is the starting height of the object. Since the bullet is fired level to the ground, it's vertical velocity is 0, and c - h(t)equals -0.1 meters, so:
-0.1=-4.9*t^2
t^2=0.02040816326530612244897959183673
t=1/7 sec
200 m. in 1/7 sec=1400 m/sec as the velocity of the bullet
B.)8 cm=.08 m.
.08 m. in 1/7 sec=.56 m/s as the velocity of the wind
☺☺☺☺